OpenStudy (nory):

How do I solve a logarithmic equation? For example, log2 (x+4) + log2 (x) = 5.

OpenStudy (anonymous):

is that log base 2 ???

OpenStudy (anonymous):

or log base 10 ???

OpenStudy (nory):

Yes. I haven't gotten the hang of the controls yet.

OpenStudy (anonymous):

???

OpenStudy (anonymous):

i will bet base 2

OpenStudy (anonymous):

is this the problem: \(\large log_2(x+4)+log_2(x)=5 \) or is it: \(\large log2(x+4)+log2(x)=5 \)

OpenStudy (nory):

It is the former.

OpenStudy (anonymous):

@satellite73 ... ding, ding, ding, ding!!!

OpenStudy (anonymous):

ok... do you know this rule for logs: \(\large log_bM+log_bN=log_b(M\cdot N) \) ???

OpenStudy (nory):

Yes, I do. So I use that rule and then what do I do? I think I know the answer but I could use a refresher.

OpenStudy (anonymous):

so you can apply that to the left side of the equation to get : \(\huge log_2[x(x+4)]=5 \) we good so far?

OpenStudy (nory):

Yes. I understand. So then I expand x(x+4)?

OpenStudy (anonymous):

yes... but you can do that later...

OpenStudy (anonymous):

do you know how to write this logarithmic equation in exponential form: \(\large log_{10}(1000)=3 \) ???

OpenStudy (nory):

10 ^ 3 = 1000. Yes. So I write the equation in exponential form?

OpenStudy (anonymous):

yes.... :)

OpenStudy (anonymous):

wat equation u got now?